a.      shows a bead sliding on a wire, How large must height h1 be if the bead, starting at rest A, is to have a velocity of 200cm/s at point B? Ignore friction



1
Expert's answer
2021-05-10T13:57:59-0400

When a bead has dropped a distance h it has lost potential energy = mgh, where m is the mass of the object, and g is the acceleration due to gravity, 9.8 "m\/s^2" . This potential energy will become kinetic energy "\frac{1}{2}mv^2" .

"v = 200 \;cm/s = 2.00 \;m/s \\

m \times 9.8 \times h = \frac{1}{2} m \times 2.0^2 = m \times \frac{4.00}{ 2} = m \times 2 \\

m \times 9.8 \times h = m \times 2 \\

9.8 \times h = 2 \\

h = \frac{2}{9.8} = 0.204 \;m = 20.4 \;cm"

Answer: 20.4 cm